# spacing=nonfrench

### v1

@@ Abstract
Recent innovations in generative large language models (LLMs) have made their applications and use-cases ubiquitous. This has led to large-scale deployments of these models, using complex, expensive, and power-hungry AI accelerators, most commonly GPUs. These developments make LLM inference efficiency an important challenge. Based on our extensive characterization, we find that there are two main phases during an LLM inference request: a compute-intensive prompt computation, and a memory-intensive token generation, each with distinct latency, throughput, memory, and power characteristics. Despite state-of-the-art batching and scheduling, the token generation phase underutilizes compute resources. Specifically, unlike compute-intensive prompt computation phases, token generation phases do not require the compute capability of the latest GPUs, and can be run with lower power and cost.
With Splitwise, we propose splitting the two phases of a LLM inference request on to separate machines. This allows us to use hardware that is well-suited for each phase, and provision resources independently per phase. However, splitting an inference request across machines requires state transfer from the machine running prompt computation over to the machine generating tokens. We implement and optimize this state transfer using the fast back-plane interconnects available in today’s GPU clusters.
We use the Splitwise technique to design LLM inference clusters using the same or different types of machines for the prompt computation and token generation phases. Our clusters are optimized for three key objectives: throughput, cost, and power. In particular, we show that we can achieve 1.4×1.4$\times$1.4 × higher throughput at 20% lower cost than current designs. Alternatively, we can achieve 2.35×2.35$\times$2.35 × more throughput with the same cost and power budgets.
@@ I Introduction
Generative Large Language Models (LLMs) have seen a lot of progress in response quality and accuracy recently [15, 43]. This has led to a wide adoption of LLMs for various use-cases [4, 17]. Most modern LLMs are based on transformers [46, 45], and share very similar characteristics [36]. Most of these models are large, and run on expensive and power-hungry GPUs [14]. The sudden large-scale deployment of LLMs has led to a world-wide GPUs capacity crunch [12].
While it is important to train these LLMs efficiently, the bulk of the datacenters and machines are being used for inference instead. The reason is the vast number of use-cases that leverage LLMs. Furthermore, the cost of training these models is very high and requires dedicated super-computers [35, 31]. A large number of inferences is the way to amortize/offset the high training costs. LLM inference jobs, although orders of magnitude smaller than training, are still expensive given the compute involved. The model size (the number of parameters in transformers models) has grown steadily, from the early BERT models [24] having 340 million parameters to GPT3 [22] with 175 billion parameters, and GPT4 rumored to have even more.
@@
TFLOPs 19.5 66.9 3.43$\times$
HBM capacity 80GB 80GB 1.00$\times$
HBM bandwidth 2039GBps 3352GBps 1.64$\times$
Power 400W 700W 1.75$\times$
NVLink 50 Gbps 100Gbps 2.00$\times$
Infiniband 200GBps 400GBps 2.00$\times$
Cost per machine [3] \$17.6/hr \$38/hr 2.16$\times$
TABLE I: NVIDIA A100 vs. H100 relevant specifications.
@@
Generative LLM inference in all the current models for a single request consists of several forward passes of the model, as the output tokens are generated one by one. This inherently has two contrasting phases of computation. First, the prompt computation phase where all the tokens in the input prompt run through the forward pass of the model in parallel, to generate the first output token. This phase tends to be computationally intensive and requires the high FLOPs (floating point operations per second) of the latest GPUs today. Second, the token generation phase which tends to be more serialized in nature, as each token is generated based on the forward pass of the last token and all the cached context from previous tokens in the sequence. Given the lack of parallelism in this computation, this phase tends to be more memory bandwidth and capacity bound, despite state-of-the-art batching. Running both phases on the same machine leads to inconsistent end-to-end latency. Due to these challenges, services need to over-provision these expensive GPUs to meet tight inference service level objectives (SLOs). On the other hand, cloud service providers (CSPs) are building many new datacenters for GPU expansions, and running into a power wall [16].
@@
Meanwhile, the industry continues to release more and more computationally able GPUs, each much more power-hungry and expensive than the previous one. However, as shown in Table I, the high-bandwidth memory (HBM) capacity and bandwidth on these GPUs has not scaled at the same rate recently. The latest NVIDIA H100 GPUs have 3.43×3.43$\times$3.43 × more compute and 1.75×1.75$\times$1.75 × more power as compared to their predecessor A100 GPUs. But the memory bandwidth increase was limited to 1.6×1.6$\times$1.6 ×, with no increase in memory capacity.
@@
Our work. Given the clear phases in the computation of an inference request for generative LLMs, we propose separating the prompt computation and the token generation phases onto separate machines. This increases hardware utilization and therefore the overall efficiency of the system. It also enables using different, better-provisioned hardware for each phase To realize such a setup, the cached context from the prompt computation will have to be communicated over from the prompt machine to the token machine with low latency. We implement these transfers in an optimized way using available bandwidths, which allows us to increase efficiency without any perceived performance loss.
@@
With Splitwise, we design clusters optimized for cost, throughput, and power, based on production traces of LLM inference requests. Given the aforementioned diverging memory and compute expansions over the generations of GPUs, we also evaluate different GPUs and power caps for the different inference phases. This allows us to target better performance per dollar (Perf/\$) for users, and better performance per watt (Perf/W) for CSPs. Additionally, users are also to target older GPUs, which are likely more readily available to them.
We show that with Splitwise-based LLM inference clusters, we can achieve 1.4$\times$ higher throughput at 20% lower cost. Alternatively, we can achieve 2.35$\times$ more throughput with the same cost and power budgets.
Summary. In this paper, we make the following contributions:
@@
    An extensive characterization of the differences in the execution and utilization patterns of the prompt and token generation phases using production traces on the two latest flagship GPUs from NVIDIA: A100 and H100.
@@
    Splitwise, our technique for optimized utilization of available hardware by separating the prompt computation and token generation phases onto separate machines.
@@
    Designing homogeneous and heterogeneous clusters with Splitwise to optimize cost, throughput, and power.
@@
    And finally an evaluation of the systems designed with Splitwise, using production traces.
@@ II-A Large Language Models
Modern LLMs are based on transformers. Transformer models use attention [45] and multi-layer-perceptron layers to understand the inputs and generate an output, respectively. Transformer-based LLMs can consist of encoder-only [24, 29], decoder-only [40, 41, 43] or encoder-decoder [42] models. Generative LLMs, the focus of this paper, are usually either decoder-only or encoder-decoder models.
@@ II-B Generative LLM inference phases
Figure 1 shows an example of generative LLM inference. Once the prompt query is received, all the input tokens are computed in parallel, in a single iteration to generate the first token. We term this as the prompt processing phase. The context generated from the attention layers during the prompt computation is saved in the key-value (KV) cache, since it is needed for all the future token generation iterations. After the first token is generated, the following tokens only use the last generated token and the KV-cache as inputs to the forward pass of the model. This makes the subsequent token generation more memory bandwidth and capacity intensive than the computationally heavy prompt phase.
In this paper, we separate the prompt and token phases of a generative LLM inference onto separate machines.
@@ II-D Batching of requests
Requests reaching an inference machine can be batched for higher throughput. Several previous works have explored batching [48, 19]. Figure 2 shows the timelines for inference with three common batching mechanisms. The default mechanism only batches at the request-level (Figure 2(a)). In this case, ready requests are batched together, but all the forward passes for these requests are completed before any other requests are run. Since requests can have long token generation phases, this can lead to long wait times for requests arriving in between, causing high TTFT and high E2E latencies. An optimization is continuous batching [48] (Figure 2(b)). In this case, the scheduling decisions are made before each forward pass. However, any given batch comprises either purely of prompt phase, or only token phase. Prompt phase is considered more important since it impacts the prompt phase (i.e., TTFT). Hence, a waiting prompt can preempt a token phase. Although this leads to shorter TTFT, it can increase the tail for TBT, and therefore E2E by a lot. Finally, there is mixed batching (Figure 2(c)) [19]. With this batching, the scheduling decisions are made at each forward pass, and the prompt and token phases can run together. This reduces the impact on TBT, but does not eliminate it, since token phases scheduled with prompt phases will experience a longer runtime. In the rest of the paper, we use mixed batching unless stated otherwise.
@@ II-E Model parallelism
Given the increasing model sizes, model parallelism is no longer just applicable to training, but also inference. Model parallelism can be used to divide a model onto multiple GPUs, and even multiple machines. There are two types of model parallelism used in inference: pipeline and tensor. Pipeline parallelism (PP) divides the layers of the model among the GPUs, while keeping all the operators and tensors within a layer on the same GPU. Tensor parallelism (TP) on the other hand, divides the tensor across the GPUs, while replicating all the layers on each GPU. Pipeline parallelism requires lower communication across the participating GPUs, while tensor parallelism requires high bandwidth communication for each layer. In general, tensor parallelism is known to be better performing for GPUs within the same machine, connected with very high bandwidth interconnect. In the rest of the paper, we use tensor parallelism across 8 GPUs for the best latency.
@@ II-F GPU clusters and interconnects
With the rise of LLM use-cases in the past years, several CSPs have expanded the GPU-based offerings, leading to large GPU cluster deployments [31, 32, 3]. Each machine in these AI clusters generally comprises 8 flagship NVIDIA GPUs (A100 or H100). Each GPUs is connected to all the other GPUs in the cluster with high bandwidth interconnects like Mellanox InfiniBand [11, 8], forming a high bandwidth data plane. The InfiniBand bandwidth offered in the cloud today ranges from 25GBps per GPU to 50GBps per GPU [8, 5]. The default optimized communication library across GPUs over NVLink or InfiniBand is NCCL [13].
@@ III Characterization
In this section, we explore the performance and utilization characteristics of LLM inference, and draw key insights to guide the design of Splitwise.
Production traces. We use production traces taken from a large cloud provider, Azure. The traces are 20 minutes long and include the arrival time, input size (number of prompt tokens), and output size (number of output tokens). We use traces representing the most common scenarios in LLM inference today: coding and conversation. Since LLM output quality has no impact on the issues we explore in this work, for each request in the trace, we use randomized text inputs that match the number of input tokens, and force the LLM to generate the right number of tokens as per the trace. For this characterization, we do not reuse any KV-cache between requests to emulate a cloud service with security guarantees. We do account for KV-cache transfers in the evaluation.
Models. Table III shows the models that we evaluate. Both BLOOM [41] and Llama2 [43] are open source models, considered to be state of the art in the generative LLM community. Both models are decoder-only transformer-based models. We use the version of each model with the most parameters, since these versions are the most representative of production-class accuracy. Unless stated otherwise, we run BLOOM-176B and Llama2-70B on vLLM [27] on a machine with 8 H100 [14] GPUs.
@@ III-A Number of prompt and generated tokens
To better understand these traces, we look at the distribution in the number of prompt input and generated output tokens. Figure 2(a) shows the distribution of number of prompt tokens. Since the coding LLM inference service is generally used to generate completions, as the user is writing code, the input prompt to this inference can include large chunks of the code written so far. This shows in the distribution, as the median prompt size is 1500 tokens. On the other hand, the conversation service has a wider range of input prompt tokens, as it depends on the user. The median number of prompt tokens for this trace is 1020 tokens.
@@
Figure 2(b) shows the distribution in the number of generated tokens. Since the common case for the coding service would only need to generate the next few words in the program as the user is typing, the median number of output token is 13 tokens. On the other hand, the conversation service has an almost bimodal distribution, with a median of 129 tokens generated.
@@
Figure 4: Cumulative distribution of time spent with various batched tokens.
Refer to caption (a) Refer to caption (b) Refer to caption (c)
Figure 4: Cumulative distribution of time spent with various batched tokens.
Figure 5: Latencies for BLOOM-176B and Llama2-70B on DGX-H100.
@@ III-B Batch utilization
To understand how much these requests can be batched, we measure how often machines run at a given batch size. We use mixed continuous batching as shown in Figure 2. To fit into a single machine, we run a scaled-down version of the coding and conversation traces with 2 requests per second. We choose the maximum request rate at which the service-level-objectives (SLOs in Table VI) can still be met.
Figure 5 shows the distribution of the time spent by the machine running various numbers of active tokens in a batch. Note that if a prompt of 100 tokens is running in its prompt phase, we count the active tokens as 100. However, once the request is in the token phase, we count it as one active token, since the tokens are generated serially, one at a time (assuming a beam search of one [27]). We find that most of the time (60–70%) for conversation is spent running only 20 tokens or less. Since the coding service has very few output tokens, it experiences even worse batching in the token phase and runs with a single token for more than 20% of the time. Both the LLMs show very similar trends.
Insight II: Most of the time with mixed continuous batching is spent in the token phase, with very few active tokens batched.
@@ III-C Latency
TTFT. Figure 4(a) shows the impact of the number of prompt tokens on TTFT. The range of sizes was chosen based on the coding and conversation traces. We find that TTFT grows almost linearly as the prompt size increases, for both the models. Prompt phase being computationally bound utilizes the GPU-compute well and causes this pattern.
TBT. Figure 4(b) shows the impact of forcefully batching the output tokens of different requests together on TBT. We observe that there is very little impact on TBT as the batch size grows, since the token phase is memory bound. With a batch size of 64, there is only <2×<2$\times$< 2 × impact on TBT.
E2E. Figure 4(c) shows various percentiles of E2E latency for coding and conversation traces running on the two models, without any batching. The variability between the request input and output sizes is apparent. Furthermore, we see that most of the E2E time is spent running the token phase of the inference. This is true even for the coding trace, where the prompt size is large, and the generated tokens are few. This further explains the findings in Figure 5. In fact, by our analysis, we find that for BLOOM-176B, a prompt phase with 1500 input tokens takes the same time as token phase with 6 output tokens.
@@
Figure 6: Impact of batching on the throughput for the two LLMs.

### v2

@@ Abstract
Generative large language model (LLM) applications are growing rapidly, leading to large-scale deployments of expensive and power-hungry GPUs. Our characterization of LLM inference shows that each inference request undergoes two phases: a compute-intensive prompt computation phase and a memory-intensive token generation phase, each with distinct latency, throughput, memory, and power characteristics. Despite state-of-the-art batching and scheduling, the token generation phase underutilizes compute resources. Unlike prompt computation, token generation does not need the compute capability of the latest GPUs and can be run with lower power and cost.
Based on these insights, we propose Splitwise, a model deployment and scheduling technique that splits the two phases of LLM inference requests on to separate machines. Splitwise enables phase-specific resource management using hardware that is well suited for each phase. Request state is transferred efficiently between machines using optimized network libraries on the fast back-plane interconnects available in today’s GPU clusters. Using Splitwise, we design homogeneous and heterogeneous LLM inference clusters optimized for throughput, cost, and power. Compared to current designs, Splitwise clusters achieve up to $1.4\times$1.4 × higher throughput at 20% lower cost. Alternatively, they can deliver $2.35\times$2.35 × more throughput under the same power and cost budgets.
@@ I Introduction
Recent advancements in generative large language models (LLMs) have significantly improved their response quality and accuracy [18, 71]. These trends have led to the widespread adoption of LLMs across various domains [6, 21]. Most modern LLMs are built using the transformer architecture [78, 77] and exhibit similar characteristics [63]. Transformer model sizes have grown steadily, from the early BERT models [36] having 340 million parameters, to GPT-3 [28] with a staggering 175 billion parameters, and GPT-4 rumored to have even more.
LLMs typically run on expensive and power-hungry GPUs [16]. The sudden and large-scale deployment of LLMs has led to a worldwide GPU capacity crunch [14]. The computational demand for LLM inference far exceeds that of training due to the vast number of applications leveraging LLMs. Furthermore, since training LLMs requires expensive and dedicated supercomputers [60, 56], a large number of inferences are necessary to amortize the high training costs. LLM inference jobs, although orders of magnitude smaller than training, are still expensive given the compute involved. 11footnotetext: Work partly done as an intern at Microsoft.
@@
[…]
TABLE I: NVIDIA A100 vs. H100 specifications.
@@
Generative LLM inference for a single request consists of several forward passes through the model, since the output tokens are generated one by one. This inherently has two contrasting phases of computation. First, the prompt computation phase, in which all the input prompt tokens run through the forward pass of the model in parallel to generate the first output token. This phase tends to be computationally intensive and requires the high FLOPs (floating point operations per second) of the latest GPUs today. Second, the token generation phase, in which subsequent output tokens are generated sequentially based on the forward pass of the last token and all the cached context from previous tokens in the sequence. Given the lack of compute parallelism, this phase tends to be more memory bandwidth and capacity bound, despite state-of-the-art batching. Running both phases on the same machine often leads to inconsistent end-to-end latencies due to the arbitrary batching of prompt and token phases. Due to these challenges, services need to over-provision expensive GPUs to meet tight inference service level objectives (SLOs) for interactive applications. At the same time, cloud service providers (CSPs) are having to build a lot of new datacenters to meet the GPU demand, and are running into a power wall [19].
@@
The industry continues to release new computationally powerful GPUs, each much more power hungry and expensive than the last. However, as shown in Table I, the high-bandwidth memory (HBM) capacity and bandwidth on these GPUs has not scaled at the same rate recently. The latest NVIDIA H100 GPUs have $3.43\times$3.43 × more compute and $1.75\times$1.75 × more power compared to their predecessor A100 GPUs. However, their memory bandwidth only grew by $1.6\times$1.6 ×, with no increase in memory capacity.
@@
Our work. Given the distinct properties of prompt computation and token generation phases, we propose splitting the inference request and running them on separate machines. Doing so allows us to separately manage hardware resources for each phase, thereby increasing the GPU utilization and the overall efficiency of the system. It also enables using different, better-suited hardware for each phase. To realize such a setup, the cached context from the prompt computation needs to be communicated over from the prompt processing machine to the token generation machine at low latency. We implement these transfers in an optimized manner over the back-end Infiniband interconnects avaialble in datacenters today, allowing us to increase efficiency without any perceived performance loss.
@@
With Splitwise, we design clusters optimized for cost, throughput, and power, using production traces of LLM inference requests [4]. Given the diverging memory and compute scaling rates across GPU generations, we also evaluate different GPUs and power caps for the different inference phases. This allows us to target better performance per dollar (Perf/\$) for users, and better performance per watt (Perf/W) for CSPs. Additionally, users can target older GPUs, which are likely more readily available to them.
[…]
Summary. We make the following contributions:
@@
    An extensive characterization of the differences in the execution and utilization patterns of the prompt and token generation phases in LLM inference on the NVIDIA A100 and H100 GPUs using production traces.
@@
    Splitwise, our technique for optimized utilization of available hardware, which splits the prompt computation and token generation phases onto separate machines.
@@
    A design exploration of homogeneous and heterogeneous cluster deployments with Splitwise to optimize the overall cost, request throughput, and provisioned power.
@@
    An evaluation of the systems designed with Splitwise using production traces.
@@ II-A Large Language Models
Modern LLMs are based on transformers. Transformer models use attention [77] and multi-layer-perceptron layers to understand the inputs and generate an output, respectively. Transformer-based LLMs include encoder-only [36, 54], decoder-only [67, 69, 71], and encoder-decoder [70] models. Generative LLMs, the focus of this paper, are usually either decoder-only, or encoder-decoder models.
@@ II-B Generative LLM inference phases
Figure 1 shows an example of generative LLM inference. Once the prompt query is received, all the input tokens are computed in parallel, within a single iteration, to generate the first token. We call this the prompt processing phase. The context generated from the attention layers during the prompt computation is saved in the key-value (KV) cache, since it is needed for all the future token generation iterations. After the first token is generated, the following tokens only use the last generated token and the KV-cache as inputs to the forward pass of the model. This makes the subsequent token generation more memory bandwidth and capacity intensive than the computationally heavy prompt phase.
@@ II-D Batching of requests
Inference requests can be batched together for higher throughput. Several prior works have explored batching [81, 23]. Figure 2 shows the timelines for inference with three common batching mechanisms. The default mechanism only batches at the request-level (Figure 2(a)). In this case, ready requests are batched together, but all the forward passes for these requests are completed before any other requests are run. Since requests can have long token generation phases, this can lead to long wait times for requests arriving in between, causing high TTFT and high E2E latencies. An optimization is continuous batching [81] (Figure 2(b)). In this case, scheduling decisions are made before each forward pass of the model. However, any given batch comprises either only of requests in their prompt phase or only requests in token phase. Prompt phase is considered more important since it impacts TTFT. Hence, a waiting prompt can preempt a token phase. Although this leads to shorter TTFT, it can substantially increase the tail for TBT, and therefore the E2E latency. Finally, there is mixed batching (Figure 2(c)) [23]. With this batching, the scheduling decisions are made at each forward pass, and the prompt and token phases can run together. This reduces the impact on TBT, but does not eliminate it, since token phases scheduled with prompt phases will experience a longer runtime. In the rest of the paper, we use mixed batching unless stated otherwise.
@@ II-E Model parallelism
Model parallelism can be used to divide a model onto multiple GPUs, and even multiple machines, for higher efficiency and memory capacity. LLM inference typically uses pipeline and tensor parallelism. Pipeline parallelism (PP) divides the layers of the model among the GPUs, while keeping all the operators and tensors within a layer on the same GPU. Tensor parallelism (TP) divides the tensor across the GPUs, while replicating all the layers on each GPU. Pipeline parallelism requires lower communication across the participating GPUs, while tensor parallelism requires high bandwidth communication for each layer. In general, tensor parallelism performs better for GPUs within the same machine, connected with high bandwidth interconnects like e.g. NVLink [15]. In the rest of the paper, we use tensor parallelism across 8 GPUs for the best latency.
@@ II-F GPU clusters and interconnects
With the recent rise of LLM use cases, several cloud service providers have expanded the GPU-based offerings, leading to large GPU cluster deployments [56, 57, 5]. Each machine in these AI clusters is generally comprised of 8 flagship NVIDIA GPUs (A100 or H100). Each GPU is connected to all the other GPUs in the cluster with a high bandwidth Mellanox InfiniBand interconnect [13, 10], forming a high bandwidth data plane network. The InfiniBand bandwidth offered in the cloud today ranges from 25 to 50GBps per GPU pair [10, 7].
@@ III Characterization
In this section, we explore the performance and utilization characteristics of LLM inference and draw key insights to guide the design of Splitwise.
Production traces. We use production traces taken from two Azure LLM inference services on November 11th 2023. Our traces represent the most common scenarios in LLM inference today: coding and conversation. We have released a subset of our traces at https://github.com/Azure/AzurePublicDataset [4]. The traces we use for characterization are 20 minutes long and include the arrival time, input size (number of prompt tokens), and output size (number of output tokens). Due to customer privacy requirements (e.g., GDPR), we do not have visibility into the content of the prompts. We instead use the production traces to guide the input and output sizes, where we send the input prompt with the required number of tokens, and force the model to generate the corresponding number of output tokens for each request. Note that the text of the inputs prompts does not impact the performance metrics that we benchmark, since they depend only on the input and output sizes. For this characterization, we do not reuse the KV-cache between requests to emulate a cloud service with security guarantees.
Models. Table III shows the models that we evaluate. Both BLOOM [69] and Llama2 [71] are state-of-the-art open source LLMs. Both models are decoder-only, transformer-based models. We use the version of each model with the most parameters, since these versions are the most representative for production-class accuracy. Unless stated otherwise, we run BLOOM-176B and Llama-70B on vLLM [51] on a machine with 8 H100 [16] GPUs.
@@ III-A Number of prompt and generated tokens
To better understand our traces, we examine the distribution of the number of prompt input and generated output tokens. Figure 3(a) shows the distribution of number of prompt tokens. Since the coding LLM inference service is generally used to generate completions as the user is writing code, its input prompt can include large chunks of the code written so far. Thus, it has a large median prompt size of 1500 tokens. On the other hand, the conversation service has a wider range of input prompt tokens since it depends on the user. The median number of prompt tokens for this trace is 1020 tokens.
@@
Figure 3(b) shows the distribution of the number of generated tokens. Since the coding service typically only generates the next few words in the program as the user types, the median number of output token is 13 tokens. On the other hand, the conversation service has an almost bimodal distribution, with a median of 129 tokens generated.
@@ III-B Batch utilization
Figure 4: Cumulative distribution of time spent with various active batched tokens.
To understand how much can these requests be batched, we measure how often machines run at a given batch size. We use mixed continuous batching as shown in Figure 2. To fit into a single machine, we run a scaled-down version of the coding and conversation traces with 2 requests per second.
Figure 4 shows the distribution of the time spent by the machine running various number of active tokens in a batch. Note that if a prompt of 100 tokens is running in its prompt phase, we count the active tokens as 100. However, once the request is in the token phase, we count it as one active token, since the tokens are generated one at a time (assuming a beam search size of one [51]). We find that most of the time (60–70%) for conversation is spent running only 20 tokens or fewer. Since the coding service has very few output tokens, it experiences even worse batching in the token phase and runs with a single token for more than 20% of the time. Both the LLMs show very similar trends.
Insight II: Mixed continuous batching spends most of the time with very few active tokens batched.
Figure 5: TTFT, TBT, and E2E for BLOOM-176B and Llama-70B on DGX-H100.
@@ III-C Latency
TTFT. Figure 5(a) shows the impact of the number of prompt tokens on TTFT. The range of sizes was chosen based on the coding and conversation traces. We find that TTFT for both models grows almost linearly as the prompt size increases. This behavior is due to the prompt phase having high GPU utilization and being computationally bound.
TBT. Figure 5(b) shows the impact of forcefully batching the output tokens of different requests together on the TBT. We observe very little impact on TBT as the batch size grows. With a batch size of 64, there is only $2\times$2 × impact on TBT.
E2E. Figure 5(c) shows various percentiles of E2E latency for both models, with no batching. The variability between the request input and output sizes is apparent. Furthermore, we see that most of the E2E time is spent running the token phase. This holds true even for the coding trace, where prompt sizes are large and generated tokens few. In fact, we find that for BLOOM-176B, a prompt phase with 1500 input tokens takes the same time as token phase with only 6 output tokens.
@@
Figure 6: Impact of batching on the throughput for the 2 LLMs.